how do i solve 2x-3y=-25
3x+y=1 by substitution

Answer by stanbon(75887)

solve 2x-3y=-25
3x+y=1 by substitution
------
Here's one way:
y = (2x+25)/3
y = (1-3x)
-----
Substitute y for y:
1-3x = (2x+25)/3
3-9x = 2x+25
11x = -22
x = -2
----
Solve for "y":
y = 1-3x
y = 1-3*-2
y = 7
==================
Cheers,
Stan H.
================